# The Identification Power of

### v1

@@ Appendix C Proofs
We begin with showing condition (i). Construction (29) implies that $Y=DY_{1}+(1-D)Y_{0}$ a.s. Hence $F$ satisfies Assumption 2.1. Similarly, construction (27) leads to $\mathbb{P}_{F}(S=D|G=\mathrm{exp},X)=1$ a.s., meaning that $F$ satisfies Assumption 2.2.
Regarding Assumption 2.3, by construction (26), $F_{Y_{1}Y_{0}|DGX}(\cdot,\cdot|d,\mathrm{exp},x)$ does not depend on $d$; i.e., $F_{Y_{1}Y_{0}|DGX}(\cdot,\cdot|1,\mathrm{exp},x)=F_{Y_{1}Y_{0}|DGX}(\cdot,\cdot|0,\mathrm{exp},x)$. Hence, $F$ satisfies Assumption 2.3.
As for Assumption 2.4, by construction (26), $F_{Y_{1}Y_{0}|DGX}(\cdot,\cdot|d,g,x)$ does not depend on $d$ and $g$. We hence have $F_{Y_{1}Y_{0}|GX}(\cdot,\cdot|\mathrm{exp},x)=F_{Y_{1}Y_{0}|GX}(\cdot,\cdot|\mathrm{obs},x)$. Note also that equation (27) implies that
$F_{S|Y_{1}Y_{0}DGX}(s|Y_{1},Y_{0},D,\mathrm{obs},X)$ $=F_{D|Y_{1}Y_{0}GX}(s|Y_{1},Y_{0},\mathrm{obs},X).$
Combining this with equation (28), we have
$F_{S|Y_{1}Y_{0}DGX}(s|Y_{1},Y_{0},d,g,X)$ $=F_{D|Y_{1}Y_{0}GX}(s|Y_{1},Y_{0},\mathrm{obs},X),$
for any $d$ and $g$. Hence $F_{S|Y_{1}Y_{0}DGX}$ does not depend on $d$ and $g$, leading to
$F_{S|Y_{1}Y_{0}GX}(s|Y_{1},Y_{0},g,X)=F_{S|Y_{1}Y_{0}X}(s|Y_{1},Y_{0},X).$
We then have
$F_{Y_{1}Y_{0}S|GX}(y_{1},y_{0},s|g,x)$ $=F_{S|Y_{1}Y_{0}GX}(s|y_{1},y_{0},g,x)\cdot F_{Y_{1}Y_{0}|GX}(y_{1},y_{0}|g,x)$
$=F_{S|Y_{1}Y_{0}X}(s|y_{1},y_{0},x)\cdot F_{Y_{1}Y_{0}|X}(y_{1},y_{0}|x).$
This result implies that $F_{Y_{1}Y_{0}S|GX}(y_{1},y_{0},s|g,x)$ does not depend on $g$, showing that $F$ satisfies Assumption 2.4.
Assumption 2.5 is also satisfied by $F$ as $F_{DGX}=F_{DGX}^{\ast}$ (equation (25)) and $F^{\ast}$ satisfies Assumption 2.5. To sum up, we have shown that $F$ satisfies condition (i) (i.e., $F$ satisfies Assumptions 2.12.5).
We next consider condition (ii) (i.e., $F_{YDX|G}(\cdot,\cdot,\cdot|\mathrm{exp})=F_{YDX|G}^{\ast}(\cdot,\cdot,\cdot|\mathrm{exp})$ and $F_{X}=F_{X}^{\ast}$). Since $F_{Y_{1}Y_{0}|DGX}(\cdot,\cdot|d,\mathrm{exp},x)$ is constructed by a copula with conditional marginal distributions $F_{Y_{1}|X}^{\ast}(\cdot|x)$ and $F_{Y_{0}|X}^{\ast}(\cdot|x)$, by Sklar’s theorem, $F_{Y_{1}Y_{0}|DGX}(\cdot,\cdot|d,\mathrm{exp},x)$ has the conditional marginal distributions corresponding to $F_{Y_{1}|X}^{\ast}(\cdot|x)$ and $F_{Y_{0}|X}^{\ast}(\cdot|x)$. Thus we have for almost all $d$ and $x$ that
$F_{Y|DGX}(\cdot|d,\mathrm{exp},x)=F_{Y_{d}|DGX}(\cdot|d,\mathrm{exp},x)=F_{Y_{d}|X}^{\ast}(\cdot|x)=F_{Y_{|}DGX}^{\ast}(\cdot|d,\mathrm{exp},x),$ (30)
where the first and third equalities follow from Assumptions 2.1, 2.3, and 2.4. Combining this result with the fact $F_{DGX}=F^{\ast}_{DGX}$ (equation (25)) leads to $F_{YDX|G}(\cdot,\cdot,\cdot|\mathrm{exp})=F_{YDX|G}^{\ast}(\cdot,\cdot,\cdot|\mathrm{exp})$. Moreover, equation (25) leads to $F_{X}=F_{X}^{\ast}$. Hence $F$ satisfies condition (ii).
We subsequently show that $\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]=\theta$. Note first that $F_{Y_{1}Y_{0}|DGX}(\cdot,\cdot|D,\mathrm{exp},X)=F_{Y_{1}Y_{0}|X}(\cdot,\cdot|X)$ from Assumptions 2.3 and 2.4. It then follows that
@@
$=\mathbb{E}_{F_{X}}\left[\int\int\psi(y_{1},y_{0})dC\left(F_{Y_{1}|X}^{\ast}(y_{1}|X),F_{Y_{0}|X}^{\ast}(y_{0}|X)\right)\right]$
$=\mathbb{E}_{F_{X}^{\ast}}\left[\int\int\psi(y_{1},y_{0})dC\left(F_{Y_{1}|X}^{\ast}(y_{1}|X),F_{Y_{0}|X}^{\ast}(y_{0}|X)\right)\right]$
@@
where the second line follows the equation $F_{Y_{1}Y_{0}|DGX}(\cdot,\cdot|D,\mathrm{exp},X)=F_{Y_{1}Y_{0}|X}(\cdot,\cdot|X)$; the third follows from equation (26); the fourth follows from equation (25); the last follows from equation (24).
Consequently, we have shown that $F\in\mathcal{F}_{\mathrm{exp}}^{\ast}$ and $\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]=\theta$. Since this discussion holds for any $\theta\in\widetilde{\Theta}_{I}$, we obtain the result $\widetilde{\Theta}_{I}\subseteq\Theta_{I}$. ∎
@@
By Sklar’s theorem, there exists a conditional copula function $C(\cdot,\cdot|s,x)\in\mathcal{C}$ that replicates $F_{Y_{1}Y_{0}|SX}$ as
$F_{Y_{1}Y_{0}|SX}(\cdot,\cdot|s,x)=C(F_{Y_{1}|SX}(\cdot|s,x),F_{Y_{0}|SX}(\cdot|s,x)|s,x),$
for almost all $(s,x)$.
@@
where the fourth line follows from the equation $F_{Y_{d}SX}=F_{Y_{d}SX}^{\ast}$, and the last follows from the definition of $\widetilde{\Theta}_{IC}$. We thus have $\theta\in\widetilde{\Theta}_{IC}$. Since this result holds for any $\theta\in\Theta_{IC}$, we have $\Theta_{IC}\subseteq\widetilde{\Theta}_{IC}$.
We next show that $\widetilde{\Theta}_{IC}\subseteq\Theta_{IC}$. Fix any $\theta\in\widetilde{\Theta}_{IC}$. By definition of $\widetilde{\Theta}_{IC}$, there exists a conditional copula function $C(\cdot,\cdot|s,x)\in\mathcal{C}$ such that
@@
We will show that there exists a distribution function $F$ in $\mathcal{F}^{\ast}$ that replicates $\theta$ as $\theta=\mathbb{E}_{F}\left[\psi(Y_{1},Y_{0})\right]$. Specifically, we construct such a a distribution function $F$ of $(Y_{1},Y_{0},Y,D,S,G,X)$ in the following way:
$F=F_{Y|Y_{1}Y_{0}DSGX}\times F_{S|Y_{1}Y_{0}DGX}\times F_{Y_{1}Y_{0}|DGX}\times F_{DGX},$
where
@@
We will show that $F\in\mathcal{F}^{\ast}$ and $\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]=\theta$. For the former, we will specifically show that (i) $F$ satisfies Assumptions 2.12.5 (with $F^{\ast}$ replaced by $F$) and (ii) $F_{YDGX}=F_{YDGX}^{\ast}$.
We begin by verifying condition (i). Construction (37) implies that $Y=DY_{1}+(1-D)Y_{0}$ a.s. Hence $F$ satisfies Assumption 2.1. Construction (35) leads to $\mathbb{P}_{F}(S=D|G=\mathrm{obs},X)=1$ a.s., implying that $F$ satisfies Assumption 2.2.
Regarding Assumption 2.3, by construction (34), $F_{Y_{1}Y_{0}|DGX}(\cdot,\cdot|d,\mathrm{exp},x)$ does not depend on $d$. Hence, $F$ satisfies Assumption 2.3.
As for Assumption 2.4, it follows from construction (34) that for almost all $x$,
$F_{Y_{1}Y_{0}|GX}(\cdot,\cdot|\mathrm{exp},x)=\mathbb{E}_{F_{D}}[F_{Y_{1}Y_{0}|DGX}(\cdot,\cdot|D,\mathrm{exp},x)]=F_{Y_{1}Y_{0}|GX}(\cdot,\cdot|\mathrm{obs},x).$ (38)
Note that equation (35) implies that
$F_{S|Y_{1}Y_{0}DGX}(s|Y_{1},Y_{0},D,\mathrm{obs},X)$ $=F_{D|Y_{1}Y_{0}GX}(s|Y_{1},Y_{0},\mathrm{obs},X).$
Combining this with equation (36), we have
$F_{S|Y_{1}Y_{0}DGX}(s|Y_{1},Y_{0},d,g,X)$ $=F_{D|Y_{1}Y_{0}GX}(s|Y_{1},Y_{0},\mathrm{obs},X),$
for any $d$ and $g$. Hence $F_{S|Y_{1}Y_{0}DGX}$ does not depend on $d$ and $g$, leading to
$F_{S|Y_{1}Y_{0}DGX}(s|Y_{1},Y_{0},d,g,X)=F_{S|Y_{1}Y_{0}GX}(s|Y_{1},Y_{0},g,X)=F_{S|Y_{1}Y_{0}X}(s|Y_{1},Y_{0},X).$
We then have for almost all $(y_{1},y_{0},s,g,x)$,
$F_{Y_{1}Y_{0}S|GX}(y_{1},y_{0},s|g,x)$ $=F_{S|Y_{1}Y_{0}GX}(s|y_{1},y_{0},g,x)\cdot F_{Y_{1}Y_{0}|GX}(y_{1},y_{0}|g,x)$
$=F_{S|Y_{1}Y_{0}X}(s|y_{1},y_{0},x)\cdot F_{Y_{1}Y_{0}|X}(y_{1},y_{0}|x).$
Therefore, $F_{Y_{1}Y_{0}S|GX}(y_{1},y_{0},s|g,x)$ does not depend on $g$, showing that $F$ satisfies Assumption 2.4.
Assumption 2.5 is also satisfied by $F$ as $F_{DG|X}=F_{DG|X}^{\ast}$ (equation (32)) and $F^{\ast}$ satisfies Assumption 2.5. To sum up, we have shown that $F$ satisfies condition (i) (i.e., $F$ satisfies Assumptions 2.12.5 with $F^{\ast}$ replaced by $F$).
We next consider condition (ii) (i.e., $F_{YDGX}=F_{YDGX}^{\ast}$). Since $F_{Y_{1}Y_{0}|DGX}(\cdot,\cdot|d,\mathrm{obs},x)$ is constructed by a copula with conditional marginal distributions $F_{Y_{1}|DGX}^{\ast}(\cdot|d,\mathrm{obs},x)$ and $F_{Y_{0}|DGX}^{\ast}(\cdot|d,\mathrm{obs},x)$, Sklar’s theorem shows that $F_{Y_{1}Y_{0}|DGX}(\cdot,\cdot|d,\mathrm{obs},x)$ has the conditional marginal distributions corresponding to $F_{Y_{1}|DGX}^{\ast}(\cdot|d,\mathrm{obs},x)$ and $F_{Y_{0}|DGX}^{\ast}(\cdot|d,\mathrm{obs},x)$; that is, for almost all $d$ and $x$,
@@
Regarding $F_{Y_{d}|DGX}(\cdot|d,\mathrm{exp},x)$, it follows for almost all $(d,x)$ that
@@
where the first equality follows from construction (34); the second follows from equation (39); the third follows from Assumption 2.4; the last follows from Assumption 2.3.
Noting that $F_{Y|DGX}(\cdot|d,g,x)=F_{Y_{d}|DGX}(\cdot|d,g,x)$ for almost all $(d,g,x)$ from construction (37), the results (39) and (40) show that $F_{Y|DGX}(\cdot|d,g,x)=F_{Y|DGX}^{\ast}(\cdot|d,g,x)$ for almost all $(d,g,x)$. Combining this result with the fact $F_{DGX}=F^{\ast}_{DGX}$ (equation (32)) leads to $F_{YDGX}=F_{YDGX}^{\ast}$. Hence $F$ satisfies condition (ii).
@@
$=\mathbb{E}_{F_{SX}}\left[\int\int\psi(y_{1},y_{0})dC\left(F_{Y_{1}|DGX}^{\ast}(y_{1}|S,\mathrm{obs},X),F_{Y_{0}|DGX}^{\ast}(y_{0}|S,\mathrm{obs},X)\right)\right]$
$=\mathbb{E}_{F_{SX}^{\ast}}\left[\int\int\psi(y_{1},y_{0})dC\left(F_{Y_{1}|DGX}^{\ast}(y_{1}|S,\mathrm{obs},X),F_{Y_{0}|DGX}^{\ast}(y_{0}|S,\mathrm{obs},X)\right)\right]$
@@
where the second and fifth lines follow from Lemma C.1 (i), the third line follows from equation (33), and the last line follows from (54). The fourth line follows from the following:
$F_{SX}(s|x)$ $=F_{S|X}(s|x)\cdot F_{X}(x)=F_{D|GX}(s|\mathrm{obs},x)\cdot F_{X}(x)$
$=F_{D|GX}^{\ast}(s|\mathrm{obs},x)\cdot F_{X}^{\ast}(x)=F_{S|X}^{\ast}(s|x)\cdot F_{X}^{\ast}(x)$
$=F_{SX}^{\ast}(s|x),$
where the second equality follows from Lemma C.1 (ii), the third equality from (32), and the fourth from Lemma C.1 (ii). We now have shown that $\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]=\theta$.
Consequently, we have shown that $F\in\mathcal{F}^{\ast}$ and $\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]=\theta$. Since this discussion holds for any $\theta\in\Theta_{IC}$, we obtain the result $\widetilde{\Theta}_{IC}\subseteq\Theta_{IC}$. ∎
@@
We first show part (i). Under the stated conditions and the definition of $\widetilde{\Theta}_{I}$, Theorem 3.2 (i) in Fan et al. (2017) implies that $\widetilde{\Theta}_{I}=[\theta^{L},\theta^{U}]$. Proposition 3.1 then establishes that $\Theta_{I}=\widetilde{\Theta}_{I}=[\theta^{L},\theta^{U}]$.
We now turn to part (ii). Under the stated conditions and the definition of $\widetilde{\Theta}_{IC}$, applying Theorem 3.2 (i) in Fan et al. (2017) with $X$ replaced by $(S,X)$ yields $\widetilde{\Theta}_{IC}=[\theta_{L},\theta_{U}]$. Theorem 3.2 then implies that $\Theta_{IC}=\widetilde{\Theta}_{IC}=[\theta_{L},\theta_{U}]$. ∎
@@
Recall definitions of $F^{\ast,(-)}(y_{1},y_{0})$, $F^{\ast,(+)}(y_{1},y_{0})$, $F_{(-)}^{\ast}(y_{1},y_{0})$, and $F_{(+)}^{\ast}(y_{1},y_{0})$ in Section 4.1. For every $(y_{1},y_{0})$, by Jensen’s inequality, we have
@@
$\theta^{L}(x)$ $=\int_{0}^{1}\psi(F_{Y_{1}|X}^{\ast,-1}(u|x),F_{Y_{0}|X}^{\ast,-1}(1-u|x))du$
$=\int\int\psi(y_{1},y_{0})dM(F_{Y_{1}|X}^{\ast}(y_{1}|x),F_{Y_{0}|X}^{\ast}(y_{0}|x))du,$
$\theta^{U}(x)$ $=\int_{0}^{1}\psi(F_{Y_{1}|X}^{\ast,-1}(u|x),F_{Y_{0}|X}^{\ast,-1}(u|x))du$
$=\int\int\psi(y_{1},y_{0})dW(F_{Y_{1}|X}^{\ast}(y_{1}|x),F_{Y_{0}|X}^{\ast}(y_{0}|x))du,$
$\theta_{L}(s,x)$ $=\int_{0}^{1}\psi(F_{Y_{1}|SX}^{\ast,-1}(u|s,x),F_{Y_{0}|SX}^{\ast,-1}(1-u|s,x))du$
$=\int\int\psi(y_{1},y_{0})dM(F_{Y_{1}|SX}^{\ast}(y_{1}|s,x),F_{Y_{0}|SX}^{\ast}(y_{0}|s,x))du,$
@@
$\theta_{U}(s,x)$ $=\int_{0}^{1}\psi(F_{Y_{1}|SX}^{\ast,-1}(u|s,x),F_{Y_{0}|SX}^{\ast,-1}(u|s,x))du$
$=\int\int\psi(y_{1},y_{0})dW(F_{Y_{1}|SX}^{\ast}(y_{1}|s,x),F_{Y_{0}|SX}^{\ast}(y_{0}|s,x))du.$
Then, $\theta^{L}=\mathbb{E}[\theta^{L}(X)]$, $\theta^{U}=\mathbb{E}[\theta^{U}(X)]$, $\theta_{L}=\mathbb{E}[\theta_{L}(S,X)]$, and $\theta_{U}=\mathbb{E}[\theta_{U}(S,X)]$.

### v4

@@ Appendix C Proofs and a Preliminary Lemma
We begin with condition (i). From equation (31), we have $Y=DY_{1}+(1-D)Y_{0}$ a.s., and hence $F$ satisfies Assumption 2.1. Moreover, equation (29) implies $\mathbb{P}_{F}(S=D|G=\mathrm{obs},X)=1$ a.s., and hence $F$ satisfies Assumption 2.2.
Regarding Assumption 2.3, equation (28) implies that $F_{Y_{1}Y_{0}|DGX}(\cdot,\cdot|d,\mathrm{exp},x)$ does not depend on $d$ (i.e., $F_{Y_{1}Y_{0}|DGX}(\cdot,\cdot|1,\mathrm{exp},x)=F_{Y_{1}Y_{0}|DGX}(\cdot,\cdot|0,\mathrm{exp},x)$). Therefore, $F$ satisfies Assumption 2.3.
As for Assumption 2.4, by construction in equation (28), $F_{Y_{1}Y_{0}|DGX}(\cdot,\cdot|d,g,x)$ does not depend on $d$ or $g$. Hence, $F_{Y_{1}Y_{0}|GX}(\cdot,\cdot|\mathrm{exp},x)=F_{Y_{1}Y_{0}|GX}(\cdot,\cdot|\mathrm{obs},x)$; that is, $F_{Y_{1}Y_{0}|GX}$ is invariant to $G$. Moreover, equations (29) and (30) imply that
$F_{S|Y_{1}Y_{0}GX}(s|Y_{1},Y_{0},\mathrm{exp},X)=F_{S|Y_{1}Y_{0}GX}(s|Y_{1},Y_{0},\mathrm{obs},X),$
so $F_{S|Y_{1}Y_{0}GX}$ is also invariant to $G$. It then follows that, for almost all $(y_{1},y_{0},s,g,x)$,
$F_{Y_{1}Y_{0}S|GX}(y_{1},y_{0},s|g,x)$ $=\int_{-\infty}^{y_{1}}\int_{-\infty}^{y_{0}}F_{S|Y_{1}Y_{0}GX}(s|u,v,g,x)dF_{Y_{1}Y_{0}|GX}(u,v|g,x)$
$=\int_{-\infty}^{y_{1}}\int_{-\infty}^{y_{0}}F_{S|Y_{1}Y_{0}X}(s|u,v,x)dF_{Y_{1}Y_{0}|X}(u,v|x).$
Thus, $F_{Y_{1}Y_{0}S|GX}(y_{1},y_{0},s|g,x)$ does not depend on $g$, showing that $F$ satisfies Assumption 2.4.
The distribution $F$ also satisfies Assumption 2.5, since $F_{DGX}=F_{DGX}^{\ast}$ (equation (27)) and $F^{\ast}$ satisfies Assumption 2.5. Overall, we have shown that $F$ satisfies condition (i) (i.e., $F$ satisfies Assumptions 2.12.5 with $F^{\ast}$ replaced by $F$).
We next consider condition (ii), namely that $F_{YDX|G}(\cdot,\cdot,\cdot|\mathrm{exp})=F_{YDX|G}^{\ast}(\cdot,\cdot,\cdot|\mathrm{exp})$ and $F_{X}=F_{X}^{\ast}$. From equation (28), Sklar’s theorem implies that $F_{Y_{1}Y_{0}|DGX}(\cdot,\cdot|d,\mathrm{exp},x)$ has conditional marginal distributions corresponding to $F_{Y_{1}|X}^{\ast}(\cdot|x)$ and $F_{Y_{0}|X}^{\ast}(\cdot|x)$. Hence, for almost all $d$ and $x$,
$F_{Y|DGX}(\cdot|d,\mathrm{exp},x)=F_{Y_{d}|DGX}(\cdot|d,\mathrm{exp},x)=F_{Y_{d}|X}^{\ast}(\cdot|x)=F_{Y|DGX}^{\ast}(\cdot|d,\mathrm{exp},x),$
where the last equality follows from Assumptions 2.1, 2.3, and 2.4. Combining this with $F_{DGX}=F^{\ast}_{DGX}$ (equation (27)) yields $F_{YDX|G}(\cdot,\cdot,\cdot|\mathrm{exp})=F_{YDX|G}^{\ast}(\cdot,\cdot,\cdot|\mathrm{exp})$. Moreover, equation (27) leads to $F_{X}=F_{X}^{\ast}$. Hence, $F$ satisfies condition (ii).
We subsequently show that $\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]=\theta$. Note first that, by Assumptions 2.3 and 2.4, $F_{Y_{1}Y_{0}|DGX}(\cdot,\cdot|D,\mathrm{exp},X)=F_{Y_{1}Y_{0}|X}(\cdot,\cdot|X)$. It then follows that
@@
$=\mathbb{E}_{F_{X}}\left[\int\int\psi(y_{1},y_{0})dC\left(F_{Y_{1}|X}^{\ast}(y_{1}|X),F_{Y_{0}|X}^{\ast}(y_{0}|X)\middle|X\right)\right]$
$=\mathbb{E}_{F_{X}^{\ast}}\left[\int\int\psi(y_{1},y_{0})dC\left(F_{Y_{1}|X}^{\ast}(y_{1}|X),F_{Y_{0}|X}^{\ast}(y_{0}|X)\middle|X\right)\right]$
@@
where the second line uses $F_{Y_{1}Y_{0}|DGX}(\cdot,\cdot|D,\mathrm{exp},X)=F_{Y_{1}Y_{0}|X}(\cdot,\cdot|X)$, the third follows from equation (28), the fourth from equation (27), and the last from equation (26).
Consequently, we have shown that $F\in\mathcal{F}_{\mathrm{exp}}^{\ast}$ and $\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]=\theta$, implying $\theta\in\Theta_{I}$. Since this argument holds for any $\theta\in\widetilde{\Theta}_{I}$, it follows that $\widetilde{\Theta}_{I}\subseteq\Theta_{I}$. ∎
@@
By Sklar’s theorem (e.g., Nelsen, 2006, Theorem 2.3.3), there exists a conditional copula function $C(\cdot,\cdot|s,x)\in\mathcal{C}$ such that, for almost all $(s,x)$,
$F_{Y_{1}Y_{0}|SX}(\cdot,\cdot|s,x)=C(F_{Y_{1}|SX}(\cdot|s,x),F_{Y_{0}|SX}(\cdot|s,x)|s,x).$
@@
where the fourth line follows from $F_{Y_{d}SX}=F_{Y_{d}SX}^{\ast}$, and the final inclusion follows from the definition of $\widetilde{\Theta}_{IC}$. Thus, $\theta\in\widetilde{\Theta}_{IC}$. Since this holds for any $\theta\in\Theta_{IC}$, we conclude that $\Theta_{IC}\subseteq\widetilde{\Theta}_{IC}$.
We next show that $\widetilde{\Theta}_{IC}\subseteq\Theta_{IC}$. Fix any $\theta\in\widetilde{\Theta}_{IC}$. By the definition of $\widetilde{\Theta}_{IC}$, there exists a conditional copula function $C(\cdot,\cdot|s,x)\in\mathcal{C}$ such that
@@
We will show that there exists a distribution function $F\in\mathcal{F}^{\ast}$ that reproduces $\theta$ as $\theta=\mathbb{E}_{F}\left[\psi(Y_{1},Y_{0})\right]$. Specifically, we construct such a distribution function $F$ of $(Y_{1},Y_{0},Y,D,S,G,X)$ hierarchically, defined by
@@
We will show that $F\in\mathcal{F}^{\ast}$ and that $\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]=\theta$. For the former, we will specifically show that (i) $F$ satisfies Assumptions 2.12.5 (with $F^{\ast}$ replaced by $F$) and (ii) $F_{YDGX}=F_{YDGX}^{\ast}$.
We begin with condition (i). From equation (38), we have $Y=DY_{1}+(1-D)Y_{0}$ a.s. Hence, $F$ satisfies Assumption 2.1. Moreover, equation (36) implies $\mathbb{P}_{F}(S=D|G=\mathrm{obs},X)=1$ a.s. Hence, $F$ satisfies Assumption 2.2.
Regarding Assumption 2.3, equation (35) implies that $F_{Y_{1}Y_{0}|DGX}(\cdot,\cdot\mid d,\mathrm{exp},x)$ does not depend on $d$. Therefore, $F$ satisfies Assumption 2.3.
As for Assumption 2.4, equation (35) implies that, for almost all $x$,
$F_{Y_{1}Y_{0}|GX}(\cdot,\cdot|\mathrm{exp},x)$ $=\mathbb{E}_{F_{D}}[F_{Y_{1}Y_{0}|DGX}(\cdot,\cdot|D,\mathrm{exp},x)]$
$=\mathbb{E}_{F_{D}}[F_{Y_{1}Y_{0}|GX}(\cdot,\cdot|\mathrm{obs},x)]=F_{Y_{1}Y_{0}|GX}(\cdot,\cdot|\mathrm{obs},x).$
Thus, $F_{Y_{1}Y_{0}|GX}$ does not depend on $G$. Moreover, equation (37) leads to
$F_{S|Y_{1}Y_{0}GX}(s|Y_{1},Y_{0},\exp,X)$ $=F_{S|Y_{1}Y_{0}GX}(s|Y_{1},Y_{0},\mathrm{obs},X),$
Hence, for almost all $(y_{1},y_{0},s,g,x)$,
$F_{Y_{1}Y_{0}S|GX}(y_{1},y_{0},s|g,x)$ $=\int_{-\infty}^{y_{1}}\int_{-\infty}^{y_{0}}F_{S|Y_{1}Y_{0}GX}(s|u,v,g,x)dF_{Y_{1}Y_{0}|GX}(u,v|g,x)$
$=\int_{-\infty}^{y_{1}}\int_{-\infty}^{y_{0}}F_{S|Y_{1}Y_{0}X}(s|u,v,x)dF_{Y_{1}Y_{0}|X}(u,v|x).$
Therefore, $F_{Y_{1}Y_{0}S\mid GX}(y_{1},y_{0},s\mid g,x)$ does not depend on $g$, and $F$ satisfies Assumption 2.4.
The distribution $F$ also satisfies Assumption 2.5, because $F_{DG\mid X}=F_{DG\mid X}^{\ast}$ and $F^{\ast}$ satisfies Assumption 2.5. To sum up, we have shown that $F$ satisfies condition (i) (i.e., $F$ satisfies Assumptions 2.12.5 with $F^{\ast}$ replaced by $F$).
We now turn to condition (ii), namely that $F_{YDGX}=F_{YDGX}^{\ast}$. From equation (34), Sklar’s theorem implies that $F_{Y_{1}Y_{0}|DGX}(\cdot,\cdot|d,\mathrm{obs},x)$ has marginals equal to $F_{Y_{1}|DGX}^{\ast}(\cdot|d,\mathrm{obs},x)$ and $F_{Y_{0}|DGX}^{\ast}(\cdot|d,\mathrm{obs},x)$; that is, for almost all $d$ and $x$,
@@
Regarding $F_{Y_{d}|DGX}(\cdot|d,\mathrm{exp},x)$, we have for almost all $(d,x)$,
@@
where the first equality follows from equation (35), the second from equation (39), the third from Assumption 2.4, and the last from Assumption 2.3.
From equation (38), we have $F_{Y|DGX}(\cdot|d,g,x)=F_{Y_{d}|DGX}(\cdot|d,g,x)$ for almost all $(d,g,x)$. Together with results (39) and (40), this implies $F_{Y|DGX}(\cdot|d,g,x)=F_{Y|DGX}^{\ast}(\cdot|d,g,x)$ for almost all $(d,g,x)$. Combining this with $F_{DGX}=F_{DGX}^{\ast}$ (equation (33)) yields $F_{YDGX}=F_{YDGX}^{\ast}$. Hence, $F$ satisfies condition (ii).
@@
$=\mathbb{E}_{F_{SX}^{\ast}}\left[\int\int\psi(y_{1},y_{0})dC\left(F_{Y_{1}|DGX}^{\ast}(y_{1}|S,\mathrm{obs},X),F_{Y_{0}|DGX}^{\ast}(y_{0}|S,\mathrm{obs},X)\middle|S,X\right)\right]$
@@
where the second and fifth equalities follow from Lemma C.1(i), the third from equation (34), and the last from equation (32). The fourth equality follows since
$F_{SX}(s,x)$ $=\int_{(-\infty,x]}F_{S|X}(s|u)dF_{X}(u)=\int_{(-\infty,x]}F_{D|GX}(s|\mathrm{obs},u)dF_{X}(u)$
$=\int_{(-\infty,x]}F_{D|GX}^{\ast}(s|\mathrm{obs},u)dF_{X}^{\ast}(u)=\int_{(-\infty,x]}F_{S|X}^{\ast}(s|u)dF_{X}^{\ast}(u)$
$=F_{SX}^{\ast}(s,x),$
where the second equality uses Lemma C.1(ii), the third equation uses (33), and the fourth again uses Lemma C.1(ii). Thus, we have established that $\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]=\theta$.
Consequently, we have shown that $F\in\mathcal{F}^{\ast}$ and $\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]=\theta$, implying $\theta\in\Theta_{IC}$. Since this holds for any $\theta\in\widetilde{\Theta}_{IC}$, we conclude that $\widetilde{\Theta}_{IC}\subseteq\Theta_{IC}$. ∎
@@
We first prove part (i). Under the stated conditions and by the definition of $\widetilde{\Theta}_{I}$, Theorem 3.2(i) in Fan et al. (2017) implies that $\widetilde{\Theta}_{I}=[\theta^{L},\theta^{U}]$. Proposition 3.1 then yields $\Theta_{I}=\widetilde{\Theta}_{I}=[\theta^{L},\theta^{U}]$.
We now turn to part (ii). Under the stated conditions and by the definition of $\widetilde{\Theta}_{IC}$, Theorem 3.2(i) in Fan et al. (2017), applied with $X$ replaced by $(S,X)$, implies that $\widetilde{\Theta}_{IC}=[\theta_{L},\theta_{U}]$. Theorem 3.3 then yields $\Theta_{IC}=\widetilde{\Theta}_{IC}=[\theta_{L},\theta_{U}]$. ∎
@@
Recall the definitions of $F^{\ast,(-)}(y_{1},y_{0})$, $F^{\ast,(+)}(y_{1},y_{0})$, $F_{(-)}^{\ast}(y_{1},y_{0})$, and $F_{(+)}^{\ast}(y_{1},y_{0})$ in Section 4.1. For any $(y_{1},y_{0})$, by Jensen’s inequality, we obtain
@@
$\theta^{L}(x)$ $\equiv\int_{0}^{1}\psi(F_{Y_{1}|X}^{\ast,-1}(u|x),F_{Y_{0}|X}^{\ast,-1}(1-u|x))du$
$=\int\int\psi(y_{1},y_{0})dM(F_{Y_{1}|X}^{\ast}(y_{1}|x),F_{Y_{0}|X}^{\ast}(y_{0}|x)),$
$\theta^{U}(x)$ $\equiv\int_{0}^{1}\psi(F_{Y_{1}|X}^{\ast,-1}(u|x),F_{Y_{0}|X}^{\ast,-1}(u|x))du$
$=\int\int\psi(y_{1},y_{0})dW(F_{Y_{1}|X}^{\ast}(y_{1}|x),F_{Y_{0}|X}^{\ast}(y_{0}|x)),$
$\theta_{L}(s,x)$ $\equiv\int_{0}^{1}\psi(F_{Y_{1}|SX}^{\ast,-1}(u|s,x),F_{Y_{0}|SX}^{\ast,-1}(1-u|s,x))du$
$=\int\int\psi(y_{1},y_{0})dM(F_{Y_{1}|SX}^{\ast}(y_{1}|s,x),F_{Y_{0}|SX}^{\ast}(y_{0}|s,x)),$
@@
$\theta_{U}(s,x)$ $\equiv\int_{0}^{1}\psi(F_{Y_{1}|SX}^{\ast,-1}(u|s,x),F_{Y_{0}|SX}^{\ast,-1}(u|s,x))du$
$=\int\int\psi(y_{1},y_{0})dW(F_{Y_{1}|SX}^{\ast}(y_{1}|s,x),F_{Y_{0}|SX}^{\ast}(y_{0}|s,x)).$
Then $\theta^{L}=\mathbb{E}[\theta^{L}(X)]$, $\theta^{U}=\mathbb{E}[\theta^{U}(X)]$, $\theta_{L}=\mathbb{E}[\theta_{L}(S,X)]$, and $\theta_{U}=\mathbb{E}[\theta_{U}(S,X)]$.